# Selective Fine-tuning on LLM-labeled Data May Reduce Reliance on Human Annotation: A Case Study Using Schedule-of-Event Table Detection

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated their efficacy across a broad spectrum of tasks in healthcare applications. However, often LLMs need to be fine-tuned on task-specific expert-annotated data to achieve optimal performance, which can be expensive and time consuming. In this study, we fine-tune PaLM-2 (Anil et al. (2023)) with parameter efficient finetuning (PEFT) using noisy labels obtained from gemini-pro 1.0 (Google (2024)) for the detection of Schedule-of-Event (SoE) tables, which specify care plan in clinical trial protocols. We introduce a filtering mechanism to select high-confidence labels for this table classification task, thereby reducing the noise in the auto-generated labels. We show that fine-tuned PaLM-2 with those labels achieves performance that exceeds the gemini-pro 1.0 and other LLMs. Furthermore, its performance is close to a PaLM-2 fine-tuned on labels obtained from non-expert annotators. Our results show that leveraging LLM-generated labels through powerful models like gemini-pro can potentially serve as a viable strategy for improving LLM performance through fine-tuning in specialized tasks, particularly in domains where expert annotations are scarce, expensive, or time-consuming to obtain.
@@ 1 Introduction
Large Language Models (LLMs) have been found to be useful across diverse tasks like natural language understanding and generation, question-answering, summarization, programming, and creative arts (Chen et al. (2021); Radford et al. (2018, 2019); Ramesh et al. (2021)). LLMs are particularly promising in specialized fields such as healthcare, where they can significantly enhance clinical decision-making, patient care, drug discovery, and the management and utilization of medical data (Singhal et al. (2023); Ingraham et al. (2023); Tu et al. (2024a, b); Sharma et al. (2024)). However, the successful application of LLMs in specialized domains frequently depends on their ability to process and understand complex, domain-specific structured and unstructured content, which often requires fine-tuning the models with data annotated by experts (van Aken (2023)). This necessity presents considerable challenges, primarily due to the scarcity, high cost, and substantial time required to acquire expert annotations in fields like healthcare. In response to these challenges, our work investigates the use of LLM-generated labels for fine-tuning purposes, with a specific case-study on identifying Schedule-of-Event (SoE) tables in clinical trial protocols. The accurate identification of SoE tables, which outlines plan-of-care in clinical trials (also see appendix A for more details on SoE tables), plays a pivotal role in the digitization of clinical trial protocols which we briefly describe below.
@@ 1.2.2 Fine-Tuning LLM with LLM-Generated Labels
We find that enhancing the quality of the auto-generated dataset for fine-tuning PaLM-2 leads to substantial improvement in fine-tuned model performance. The PaLM-2 model trained with these subset of LLM generated labels outperforms both the baseline PaLM-2 and gemini-pro 1.0 on SoE detection task (see table 2). Remarkably, the fine-tuned PaLM-2 model achieves performance levels close to those obtained with human-annotated labels, showcasing the effectiveness of using LLM-generated labels for domain-specific tasks, particularly in settings where expert annotations are sparse.
The success of our approach on SoE table classification, a highly specialized and narrow task, underscores the broader potential of LLMs in natural language understanding. It also highlights the potential impact these models can have on streamlining and enhancing manual processes in complex domains such as clinical trials. Finally, our work shows that auto-generated labels can be effective for fine-tuning LLMs for highly specialized applications. This can offer a scalable and cost-effective alternative to traditional annotation methods relying on human annotators in specialized domain like healthcare.
Summary of Contributions
  • We present an innovative fine-tuning approach for Large Language Models (LLMs), utilizing noisy labels from another LLM for the task of table classification. This method incorporates a strategic label filtering mechanism—selecting labels for fine-tuning only when there is an agreement between dual data representations of tables—which leads to significant performance improvements compared to both the base model and the label-generating LLM and closely approaches the performance of a model fine-tuned with human annotations.
  • Our work underscores the adaptability and robust capabilities of LLMs in processing and interpreting complex, domain-specific documents, reinforcing their role as a valuable tool in automating and improving manual and error-prone workflows in specialized domains of healthcare like clinical trials.
  • We propose that LLMs can be a viable option for generating labels for fine-tuning other LLMs in specialized domains like healthcare where expert annotations are often scarce or prohibitively expensive. Auto-generation of labels can provide a scalable and economically efficient alternative to conventional human-based annotation in certain scenarios, paving the way for broader adoption and application of LLMs in data-rich, expertise-driven fields like healthcare. We also discuss broader implications and ethical considerations of generating labels using LLMs.
@@ 2.2 Model Training on Synthetic Dataset
Successes of generative models in various tasks have spurred research into leveraging these models to augment real data for model fine-tuning and validation. Besnier et al. (2019), for example, use class-conditional GAN generated image for training model for image classification tasks. He et al. (2023) study the potential of synthetic data in zero-shot and few-shot classification using CLIP model (Radford et al. (2021). Recently, researchers have also used LLMs to augment data for various classification tasks. Meng et al. (2022) use a pre-trained language model to generate samples by prompting it with real data and using the generated data for fine-tuning a BERT model. To control the quality of samples, they use log-probability of generated samples for filtering poor quality auto-generated samples. Yoo et al. (2021) use randomly sampled existing data samples to condition models to generate new samples, while using token probability corresponding to the label-classes to obtain soft probability for these generated samples. These soft probabilities for synthetic samples are used to train BERT-style models for classification. One of the recent studies by Li et al. (2023) has tried to understand when synthetic data can be helpful in successful model training. They find that synthetic data is less effective when a classification task is subjective or when a specific instance of data to be classified is subjective as measured by agreement amongst annotators.
Building on these previous research work, our study employs LLM generated labels for fine-tuning another LLM for table classification task in a highly specialized context, specifically, Schedule-of-Event table classification in clinical trial protocols. Distinct from previous research, which often relies on standard benchmarks or datasets for generating synthetic data, our work showcases a novel application of synthetic labels for fine-tuning in domains where expert annotation is expensive and challenging to obtain. Furthermore, we offer a detailed comparison between models fine-tuned on LLM-generated data versus those fine-tuned on data annotated by human experts. Our approach also introduces an innovative label filtering mechanism that utilizes dual data representations of tables for removing potentially noisy synthetic labels. Finally, our approach doesn’t require access to logits for tokens and can be applied even when working with LLMs through black-box API access.
@@ 3.2.1 Training and Test Set
Our data set consists of a total of 499 clinical trial protocols of which 91 are expert-labeled by a team of five protocol digitization specialists and are used as the test set in our experiments. These 91 test protocols have a total of 3019 tables with 411 SoE tables (13.6%) and 2608 non-SoE tables (86.4%.) These expert-digitizers are specifically trained to manually label and digitize the clinical protocol for our in-house clinical trial management system (CTMS) software and the labeling process and digitization requires significant domain knowledge, time and effort. We take the expert annotations as ground truth for all experiments. The subset of 408 protocols that don’t have any expert-labels are used for the fine-tuning tasks. Of the 408 protocols, we randomly select 300 as training set, 18 as validation set and 90 as test set for model fine-tuning. These 499 protocols in our experiments span a diverse set of clinical trials across pharmaceutical companies, academic organizations, hospitals, and government organizations.
@@ 3.3 Selective Human and LLM Annotation
We first do inference with the PaLM-2 model. On a subset of 60 protocols, we manually go through PaLM-2 model prediction with the help of experts to find specific patterns in incorrect model prediction. We find that the base PaLM-2 has a very high recall but also a very high false positive rate and often predicts trivial cases of non-SoE tables as SoE. Thus, we only obtain human and gemini-pro 1.0 annotations on tables identified as SoE by the base PaLM-2 model (around 25% of all tables.) This selective annotation allows us to keep the size of annotation tasks manageable (by reducing the task to one-fourth), while also helps us over-sample the SoE table examples for fine-tuning. Additionally, this approach allows annotators to concentrate their efforts on more ambiguous cases potentially leading to higher quality annotations since annotators can spend more time on each item.
@@
We emphasize that we do not use expert annotators directly for labeling tasks. However, the annotators do have some previous experience with annotation for SoE tables and they also have access to expert annotators for any annotation they need help with. Additionally, they can choose to not annotate a table and leave it as “Do not know”. These are later annotated by an expert. Despite access to experts, non-expert annotations can be noisy due to variation in skills among the non-expert annotators. On random overlapping sets of 50 annotation, the average inter-rater agreement among non-expert annotators is 81.2%. We note that all annotations are collected only on tables predicted as SoE by base PaLM-2 models as described previously. Thus, the inter-rate agreement is only on a portion of all tables present in the protocols that are annotated by human experts.
@@ 3.4 Experiments
For fine-tuning the PaLM-2 model, we use the 408 protocols as previously described in section 3.2.1. We fine-tune all PaLM-2 models on table annotations obtained from 300 protocols and use 18 protocols for validation and the rest 90 protocols as test set. We fine-tune two sets of models with gemini-pro 1.0 generated annotations and one model with human annotations. The first gemini-pro based fine-tuned model uses all 2800 table annotations from 300 protocols, while the second removes all “noisy labels” from fine-tuning. Specifically, for the second fine-tuning experiment with gemini-pro annotations, we remove all samples from training where the JSON and text annotations of gemini-pro are not identical. This reduces the set of table annotations to 2512 tables in the training set. All models are fine-tuned for 300 epochs with learning rate multiplier of 1, early stopping set to True, and an evaluation interval of 10 epochs with Google Cloud Vertex AI fine-tuning pipeline which uses Parameter Efficient Fine-tuning (PEFT). We track the model training through a tensorboard instance.
@@ 4 Results
We evaluate models on a comprehensive set of metrics to assess the effectiveness of each model in the context of clinical trial protocol digitization. The models are benchmarked based on recall, precision, F-1 score, and accuracy. We additionally measure model performance at various precision threshold as well as on the percentage of protocols achieving 100% recall and precision (refer appendix D), which are critical for the practical deployment of the automated digitization pipeline.
@@ 4.1 Baselines
We start with a very simple baseline of non-finetuned models–gemini-pro 1.0, GPT-4, and PaLM-2. We have summarized the results in Table 2. We notice that all baseline models achieve a very high recall. Among the models that are not fine-tuned, the inference with GPT-4 results in best performance with a precision of 78.2%, f1-score of 0.84 and an accuracy of 94.0%. The inference with the PaLM-2 base model achieves 59.8% precision, 0.71 f1-score and 87.6% accuracy. The inference with gemini-pro 1.0 results in a performance between PaLM-2 and GPT-4 with 65.7% precision, 0.76 f1-score and 90.0% accuracy. In addition to these baselines, we also use baselines with naive combinations of gemini-pro 1.0 and PaLM-2 prediction (see Appendix C for details.)
Table 2: Average Recall, Precision, F-1 Score, and Accuracy on the test set
Model Recall Precision F-1 Score Accuracy
PaLM-2 97.2% 59.8% 0.71 87.6%
GPT-4 (gpt-4-0613) 98.6% 78.2% 0.84 94.0%
Gemini Pro 1.0 99.5% 65.7% 0.76 90.0%
Fine-tuned PaLM-2 98.9% 87.3% 0.91 96.0%
(Using Human Labels)
Fine-tuned PaLM-2 100% 63.7% 0.74 88.1%
(Using ALL Gemini Labels)
Fine-tuned PaLM-2 97.7% 85.9% 0.89 95.7%
(Using Filtered Gemini Labels)

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated their efficacy across a broad spectrum of tasks in healthcare applications. However, often LLMs need to be fine-tuned on task-specific expert-annotated data to achieve optimal performance, which can be expensive and time consuming. In this study, we fine-tune PaLM-2 (Anil et al. (2023)) with parameter efficient fine-tuning (PEFT) using noisy labels obtained from gemini-pro 1.0 (Google (2024)) for the detection of Schedule-of-Event (SoE) tables, which specify care plan in clinical trial protocols. We introduce a filtering mechanism to select high-confidence labels for this table classification task, thereby reducing the noise in the auto-generated labels. We find that the fine-tuned PaLM-2 with filtered labels outperforms Gemini Pro 1.0 and other LLMs on this task and achieves performance close to PaLM-2 fine-tuned on non-expert human annotations. We also explore supplementing LLM-generated labels with a limited number of human annotations for challenging cases. Our results show that leveraging LLM-generated labels, coupled with strategic filtering and selective human annotation, can be a viable and cost-effective strategy for improving LLM performance on specialized tasks, especially in domains where expert annotations are scarce, expensive, or time-consuming to obtain.
@@ 1 Introduction
Large Language Models (LLMs) have been found to be useful across diverse tasks like natural language understanding and generation, question-answering, summarization, programming, and creative arts (Chen et al. (2021); Radford et al. (2018, 2019); Ramesh et al. (2021)). LLMs are particularly promising in specialized fields such as healthcare, where they can significantly enhance clinical decision-making, patient care, drug discovery, and the management and utilization of medical data (Singhal et al. (2023); Ingraham et al. (2023); Tu et al. (2024a, b); Sharma et al. (2024)). However, the successful application of LLMs in specialized domains frequently depends on their ability to process and understand complex, domain-specific structured and unstructured content, which often requires fine-tuning the models with data annotated by experts (van Aken (2023)). This necessity presents considerable challenges, primarily due to the scarcity, high cost, and substantial time required to acquire expert annotations in fields like healthcare. In response to these challenges, our work investigates the potential of LLM-generated labels for fine-tuning purposes, with a specific case-study on identifying Schedule-of-Event (SoE) tables in clinical trial protocols. The accurate identification of SoE tables, which outlines plan-of-care in clinical trials (also see appendix A for more details on SoE tables), plays a pivotal role in the digitization of clinical trial protocols which we briefly describe below.
@@ 1.2.2 Fine-Tuning LLM with LLM-Generated Labels
We find that enhancing the quality of the auto-generated dataset for fine-tuning PaLM-2 leads to substantial improvement in fine-tuned model performance. The PaLM-2 model trained with these subset of LLM generated labels outperforms both the baseline PaLM-2 and gemini-pro 1.0 on SoE detection task (see table 3). Remarkably, the fine-tuned PaLM-2 model achieves performance levels close to those obtained with human-annotated labels, showcasing the effectiveness of using LLM-generated labels for domain-specific tasks, particularly in settings where expert annotations are sparse. Finally, we explore a hybrid approach where a limited number of low-confidence LLM-generated labels, as identified by our filtering mechanism (i.e., instances where JSON and text-based inferences disagree), are selectively replaced with human annotations. This strategy aims to leverage the efficiency of LLM-generated labels while mitigating potential concerns regarding biases that might arise from discarding labels. Although the performance of the models fine-tuned with human and gemini annotations with the filtering mechanism is already comparable, this hybrid approach utilizing only 10% of human labeling further narrows the gap with the PaLM-2 model trained with human labels as seen in table 3.
1.3 Generalizable Insights about Machine Learning in the Context of Healthcare
Our work on SoE table classification, a highly specialized healthcare task, provides valuable insights into the potential of Large Language Models (LLMs) for advancing machine learning applications in healthcare. Our work demonstrates that LLM-generated labels, when combined with filtering and selective human validation, may offer a scalable and cost-effective alternative to traditional expert annotation for fine-tuning domain-specific models. This has significant implications for healthcare, where access to expert annotations is often limited due to cost, time constraints, and the limited availability of experts.
We summarize the key insights below:
  • LLMs may be effectively applied to highly specialized healthcare tasks and have the potential to transform manual workflows: Our results show that even for tasks that require a deep understanding of domain-specific terminology and concepts, like SoE table classification, LLMs can potentially achieve strong performance, especially after fine-tuning. This adaptability of LLMs highlights their potential to automate and streamline traditionally manual and error-prone processes in healthcare.
  • LLM-generated labels offer a promising solution for addressing annotation challenges: Leveraging LLMs to generate labels presents a scalable and cost-effective alternative to traditional manual and expert-based annotation, potentially accelerating the development of ML models in healthcare.
  • Strategic filtering and human validation of LLM generated labels are crucial for optimizing performance: Our filtering mechanism, based on dual data representations, significantly improves the quality of LLM-generated labels and subsequent fine-tuned model performance. Our hybrid labeling strategy shows that selective human annotation on difficult examples may further improve performance and bridge the gap with models fine-tuned entirely on human labels.
@@ 2.2 Model Training on Synthetic Dataset
Successes of generative models in various tasks have spurred research into leveraging these models to augment real data for model fine-tuning and validation. Besnier et al. (2019), for example, use class-conditional GAN generated image for training model for image classification tasks. He et al. (2023) study the potential of synthetic data in zero-shot and few-shot classification using CLIP model (Radford et al. (2021). Recently, researchers have also used LLMs to augment data for various classification tasks. Meng et al. (2022) use a pre-trained language model to generate samples by prompting it with real data and using the generated data for fine-tuning a BERT model. To control the quality of samples, they use log-probability of generated samples for filtering poor quality auto-generated samples. Yoo et al. (2021) use randomly sampled existing data samples to condition models to generate new samples, while using token probability corresponding to the label-classes to obtain soft probability for these generated samples. These soft probabilities for synthetic samples are used to train BERT-style models for classification. One of the recent studies by Li et al. (2023) has tried to understand when synthetic data can be helpful in successful model training. They find that synthetic data is less effective when a classification task is subjective or when a specific instance of data to be classified is subjective as measured by agreement amongst annotators. In the healthcare domain, Feder et al. (2024) have explored counterfactual data augmentation for improved LLM generalization showing promising result on text classfication tasks.
Building on these previous research, our study employs LLM generated labels for fine-tuning another LLM for table classification task in a highly specialized context, specifically, Schedule-of-Event table classification in clinical trial protocols. Distinct from previous research, which often relies on standard benchmarks or datasets for generating synthetic data, our work showcases a novel application of synthetic labels for fine-tuning in domains where expert annotation is expensive and challenging to obtain. Furthermore, we offer a detailed comparison between models fine-tuned on LLM-generated labels versus those fine-tuned on labels annotated by human experts. Our approach also introduces an innovative label filtering mechanism that utilizes dual data representations of tables for removing potentially noisy synthetic labels and allowing for selective annotation by human annotators. Finally, our approach doesn’t require access to logits for tokens and can be applied even when working with LLMs through black-box API access.
@@ 3.1 Problem Set-up
Figure 1: Schematic of the fine-tuning process. This figure illustrates our approach to fine-tuning PaLM-2 for SoE table classification using LLM-generated labels. A base PaLM-2 model is first applied to identify potential SoE tables in 408 unlabelled protocols. Only predicted SoE tables are then annotated by both Gemini Pro (using JSON and text representations) and human annotators. Our results (table 3) show that our filtering approach (where we discard Gemini-pro labels with different inferences for JSON and text view before fine-tuning) achieves performance close to that of fine-tuning on human labels, highlighting the potential of LLM-generated labels for specialized healthcare tasks.
@@ 3.2.1 Training and Test Set
Our data set consists of a total of 499 clinical trial protocols of which 91 are expert-labeled by a team of five protocol digitization specialists and are used as the test set in our experiments. These 91 test protocols have a total of 3019 tables with 411 SoE tables (13.6%) and 2608 non-SoE tables (86.4%.) These expert-digitizers are specifically trained to manually label and digitize the clinical protocol for our in-house clinical trial management system (CTMS) software and the labeling process and digitization requires significant domain knowledge, time and effort. We take the expert annotations as ground truth for all experiments. The subset of 408 protocols that don’t have any expert-labels are used for the fine-tuning tasks (see figure 1). Of the 408 protocols, we randomly select 300 as training set, 18 as validation set and 90 as test set for model fine-tuning task. Note that we use a separate expert-labeled test data set of 91 protocols for all model evaluations distinct from the one used during fine-tuning. These 499 protocols in our experiments span a diverse set of clinical trials across pharmaceutical companies, academic organizations, hospitals, and government organizations.
@@ 3.3 Selective Human and LLM Annotation
We show the fine-tuning process workflow in figure 1. We first do inference of all 408 unlabelled protocols with the PaLM-2 model. On a subset of 60 protocols, we manually go through PaLM-2 model prediction with the help of experts to find specific patterns in incorrect model prediction. We find that the base PaLM-2 has a very high recall but also a very high false positive rate and often predicts trivial cases of non-SoE tables as SoE. Thus, we only obtain human and gemini-pro 1.0 annotations on tables identified as SoE by the base PaLM-2 model (around 25% of all tables.) This selective annotation allows us to keep the size of annotation tasks manageable (by reducing the task to one-fourth), while also helps us over-sample the SoE table examples for fine-tuning. Additionally, this approach allows annotators to concentrate their efforts on more ambiguous cases potentially leading to higher quality annotations since annotators can spend more time on each item.
@@
We emphasize that we do not use expert annotators directly for labeling tasks. However, the annotators do have some previous experience with annotation for SoE tables and they also have access to expert annotators for any annotation they need help with. Additionally, they can choose to not annotate a table and leave it as “Do not know”. These are later annotated by an expert. Despite access to experts, non-expert annotations can be noisy due to variation in skills among the non-expert annotators. On random overlapping sets of 50 annotation, the average inter-rater agreement among non-expert annotators is 81.2%. We note that all annotations are collected only on tables predicted as SoE by base PaLM-2 models as described previously. Thus, the inter-rate agreement is only on a portion of all tables present in the protocols that are annotated by human experts. For tables with multiple annotations, we choose the first annotation collected for that table.
Finally, to get a sense of the alignment between between Gemini-pro and the non-expert labels, we looked at the overlap between these two sets of annotations. We found that 89.6% of the annotations between human labelers and gemini-pro are identical. We show the agreement matrix for human and gemini-pro labels in table 2. We can see that majority of the disagreement are cases where gemini-pro classifies a table as SoE, but human raters don’t. This is in-line with a high recall and low precision observed for gemini-pro model as seen in table 3 on the test set.
Table 2: Agreement between Gemini Pro and Non-Expert Labelers
Non-Expert Labelers
SoE Non-SoE
Gemini Pro SoE 1497 251
Non-SoE 39 1013
Overall Agreement 89.6%
@@ 3.4 Experiments
For fine-tuning the PaLM-2 model, we use the 408 protocols as previously described in section 3.2.1. We fine-tune all PaLM-2 models on table annotations obtained from 300 protocols and use 18 protocols for validation and the rest 90 protocols as test set for our fine-tuning process. We fine-tune three sets of models with gemini-pro 1.0 generated annotations and one model with human annotations. The first gemini-pro based fine-tuned model uses all 2800 table annotations from 300 protocols, while the second removes all “noisy labels” from fine-tuning. Specifically, for the second fine-tuning experiment with gemini-pro annotations, we remove all samples from training where the JSON and text annotations of gemini-pro are not identical. This reduces the set of table annotations to 2518 tables in the training set. Finally, for the third fine-tuned model, instead of discarding the 282 samples (having different text and JSON inferences), we replace them with human annotations. All models are fine-tuned for 300 epochs with learning rate multiplier of 1, early stopping set to True, and an evaluation interval of 10 epochs with Google Cloud Vertex AI fine-tuning pipeline which uses Parameter Efficient Fine-tuning (PEFT). We track the model training through a tensorboard instance.
@@ 4 Results
We evaluate models on a comprehensive set of metrics to assess the effectiveness of each model in the context of clinical trial protocol digitization. The models are benchmarked based on recall, precision, F-1 score, and accuracy. We additionally measure model performance at various precision threshold as well as on the percentage of protocols achieving 100% recall and precision (refer appendix D), which are critical for the practical deployment of the automated digitization pipeline. To estimate the uncertainty in our performance metrics, we use bootstrapping with 10,000 replications. This involves repeatedly resampling the test dataset with replacement and calculating the metrics for each resampled dataset. The 95% confidence intervals (CIs) are then derived from the distribution of these bootstrapped estimates.
@@ 4.1 Baselines
We start with a very simple baseline of non-finetuned models–gemini-pro 1.0, GPT-4, and PaLM-2. We have summarized the results in Table 3. We notice that all baseline models achieve a very high recall. Among the models that are not fine-tuned, the inference with GPT-4 results in best performance with a precision of 78.2% (73.1, 83.2), f1-score of 0.845 (0.807, 0.881) and an accuracy of 94.0% (91.2, 96.2). The inference with the PaLM-2 base model achieves 59.8% (54.7, 65.0) precision, 0.710 (0.667, 0.752) f1-score and 87.6% (84.1, 90.7) accuracy. The inference with gemini-pro 1.0 results in a performance between PaLM-2 and GPT-4 with 65.7% (60.6, 70.9) precision, 0.761 (0.719, 0.802) f1-score and 90.0% (86.8, 92.8) accuracy. In addition to these baselines, we also use baselines with naive combinations of gemini-pro 1.0 and PaLM-2 prediction (see Appendix C for details.)
Table 3: Average Recall, Precision, F-1 Score, and Accuracy on the Test Set (Values reported as mean (95% CI))
Model Type Model/Training Data Recall % Precision % F-1 Accuracy %
Non-Fine-Tuned PaLM-2 97.3 59.8 0.710 87.6
(94.9, 99.3) (54.7, 65.0) (0.667, 0.752) (84.1, 90.7)
GPT-4 (gpt-4-0613) 98.6 78.2 0.845 94.0
(97.1, 99.8) (73.1, 83.2) (0.807, 0.881) (91.2, 96.2)
Gemini Pro 1.0 99.4 65.7 0.761 90.0
(98.4, 100.0) (60.6, 70.9) (0.719, 0.802) (86.8, 92.8)
Fine-Tuned PaLM-2 Human Labels 98.9 87.3 0.908 96.0
(97.3, 100.0) (82.9, 91.4) (0.875, 0.938) (93.5, 98.0)
All Gemini labels 100 63.8 0.744 88.1
(100.0, 100.0) (58.1, 69.2) (0.698, 0.788) (84.2, 91.5)
Filtered Gemini labels 97.7 85.9 0.894 95.7
(95.5, 99.5) (81.2, 90.3) (0.858, 0.927) (93.1, 97.7)
Human + Gemini labels 98.9 86.4 0.903 95.8
(97.3, 100.0) (81.6, 90.7) (0.867, 0.934) (93.2, 97.8)